OpenStudy (anonymous):

Raul is 5 years older than twice carlos' age. The sum of their ages is 101. How old is Carlos?

OpenStudy (anonymous):

Let x be Carlos age. 2x +5 will be Raul's age. You have to solve x + 2x + 5 =101 to find x and x will be Carlos age.

OpenStudy (anonymous):

haha uh would the answerr be 30

OpenStudy (anonymous):

No 3 x = 101-5=96 How much is x?

OpenStudy (anonymous):

I dont know?91?

OpenStudy (anonymous):

Can someone please help me?

OpenStudy (anonymous):

96/3 = 32

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Oh okay thanks:)